Lars:

Is the point (2, –1) is true for the system? (Show handwritten work) 3x - y = 7 2x + 3y = 1

Lars:

I'd really appreciate an answer asap

MidniteSunny:

I don't usually help with math, but let me see what I can do. (2, -1) is the point given, ok, so I think just substitute it into both equations and see if they're true or not. 3x-y=7 3(2)-(-1)=7 6+1=7 7=7 2x+3y=1 2(2)+3(-1)=1 4+(-3)=1 1=1 So it seems that (2, -1) is true for this system of equations.

Lars:

@midnitesunny wrote:
I don't usually help with math, but let me see what I can do. (2, -1) is the point given, ok, so I think just substitute it into both equations and see if they're true or not. 3x-y=7 3(2)-(-1)=7 6+1=7 7=7 2x+3y=1 2(2)+3(-1)=1 4+(-3)=1 1=1 So it seems that (2, -1) is true for this system of equations.
Thank you!

MidniteSunny:

ofc